OpenStudy (anonymous):

i need help please!!!! solve using factoring,complete the square,square root property or quadratic formula. A) x^2-4x+9=0 B) 1/5x-1/4x +1/3x =-17/60

OpenStudy (anonymous):

the firrst one will have complex zeros (not real ones) is that ok?

OpenStudy (anonymous):

can you explain to me please

OpenStudy (anonymous):



OpenStudy (anonymous):

is it \[x^2-4x+9=0\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

subtract 9 from both sides get \[x^2-4x=-9\] not plus 9

OpenStudy (anonymous):

omg, yes!! you are right! my mistake

OpenStudy (anonymous):

then half of 4 is 2 and 2 squared is 4 so go right to \[(x-2)^2=-9+4\] or \[(x-2)^2=-5\]

OpenStudy (anonymous):

yes! thank you, is that the final answer for part A?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are you allowed complex numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x-2)^2=-5\\ x-2=\pm\sqrt{5}i\\ x=2\pm\sqrt{5}i\] last line is the "final answer "

OpenStudy (anonymous):

does the i remain under the square root or does it come out?

OpenStudy (anonymous):

the \(i\) is the outside the square root

OpenStudy (anonymous):

\[\sqrt{-5}=\sqrt5\sqrt{-1}=\sqrt5i\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

got it, thanks so much

OpenStudy (anonymous):

could you also help me with B?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

is it \[\frac{1}{5}x-\frac{1}{4}x+\frac{1}{3}x=-\frac{17}{60}\]?

OpenStudy (anonymous):

the x's are with the denominators, i dont know if that makes a difference

OpenStudy (anonymous):

yes is certainly does!

OpenStudy (anonymous):

\[\frac{1}{5x}-\frac{1}{4x}+\frac{1}{3x}=-\frac{17}{60}\]like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply both sides by \(60x\)

OpenStudy (anonymous):

** also, the negative on the right of the equal sign was in the numerator**sorry

OpenStudy (anonymous):

that makes no difference

OpenStudy (anonymous):

\[12x-15x+10x=-17x\]

OpenStudy (anonymous):

each term is multiplied by 60x?

OpenStudy (anonymous):

yes, but there is no solution

OpenStudy (anonymous):

oh wait hold on

OpenStudy (anonymous):

i made a mistake multiply both sides by \(60x\) you get \[12-15+10=-17x\]you can solve that for \(x\)

OpenStudy (anonymous):

you get \[-7=-17x\] and so \[x=\frac{7}{17}\]

OpenStudy (anonymous):

i got that x should equal -1 :S

OpenStudy (anonymous):

1/3x * 60x is 20

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i mean oops on my part

OpenStudy (anonymous):

its okay :) thank you!! you were such a huge help

OpenStudy (anonymous):

\[12-15+20=17\] so you are right and i am wrong

OpenStudy (anonymous):

yw